(1-radical 2) (4+3 radical 5)


and

radical x+3= radical 5-x

Answer by stanbon(75887)

(1-radical 2) (4+3 radical 5)
=(1-sqrt2)*4 + (1-sqrt2)(3sqrt5)
=4-4sqrt2 + 3sqrt5--3sqrt10
-----------------
and
radical(x+3)= radical(5-x)
Square both sides to get:
x+3 = 5-x
2x=2
x=1
------
Substitute to check the answer:
sqrt(1+3) =sqrt(5-1)
True
--------
So, x=1
===================
Cheers,
Stan H.